OpenStudy (anonymous):

What is one of the products produced when Mg(NO3)2 and AlPO4 react together? Mg3(PO4)2 Al(NO3)2 Mg(PO4)2 AlNO3

OpenStudy (anonymous):

Mg3(PO4)2...i think =D

OpenStudy (anonymous):

I think A but im not positive McSpazzy will know

OpenStudy (anonymous):

either A or C..?

OpenStudy (anonymous):

\[Phosphate : PO _{4}^{-3}\]\[Magnesium : Mg ^{2+}\]

OpenStudy (anonymous):

\[Nitrate : NO _{3}^{-}\]\[Aluminium : Al ^{3+}\]

OpenStudy (anonymous):

3(2+) + 2(3-) = 0 <- A is the only valid answer. 1(3+) + 2(1-) = 2 1(2+) + 2(3-) = 4 1(3+) + 1(1-) = 2

OpenStudy (anonymous):

thanks so much guys!!

OpenStudy (anonymous):

No problem. I'll be around if you have any other problems.

OpenStudy (anonymous):

i know i will haha

OpenStudy (anonymous):

Just tag me or somethin.